W2

Most processes fail on startup (see screenshots/systprocess_fault110909.png). After a restart everything is OK.

RON TEST CCD39

time:10:48
ALL ON ,source lamp 0%
----------------------------- SUMMARY ------------------------------------
ccd39 - binning 1
Speed: 2500 kpix/sec - RON (ADU): 29.84 - RON (e-): 14.92 (estimated)
Speed: 890 kpix/sec - RON (ADU): 17.08 - RON (e-): 8.54 (estimated)
Speed: 400 kpix/sec - RON (ADU): 8.91 - RON (e-): 4.45 (estimated)
Speed: 150 kpix/sec - RON (ADU): 10.63 - RON (e-): 5.32 (estimated)

COPLEY. source lamp 0%

------------------------------ SUMMARY ------------------------------------
ccd39 - binning 1
Speed: 2500 kpix/sec - RON (ADU): 19.43 - RON (e-): 9.71 (estimated)
Speed: 890 kpix/sec - RON (ADU): 13.23 - RON (e-): 6.61 (estimated)
Speed: 400 kpix/sec - RON (ADU): 7.80 - RON (e-): 3.90 (estimated)
Speed: 150 kpix/sec - RON (ADU): 8.61 - RON (e-): 4.31 (estimated)

AdSec chiller

Secondary mirror chiller does not work, impossible to work with secondary mirror.
We wait for the service mantainance at 1:00 pm.

The systems seems that is not working and the secondary staff will try a solution which it could give more operation time

Pupil rerotator

Motor stuck during homing with a very loud sound. Decreased movement speed from "12" to "4" (in Pollux units), as done months ago during the cold test. Movement seems OK both forward and backward.

Porting the W2 software control to aoserver.

The software control of the W2 seems to be installed correctly on the aoserver.

During the software test the flower pot stopped to started to send wrong string to the serial. Controlling the PIC - doing a telnet connections - the PIC was sending wrong string. Thgis seems that the software doesn't have any control on teh devices monitored and controlled by the PIC. This problem start around 12:10 and we could get something in operation at 15:00, the problem seems something with the chip of the PIC, looking it seems that the pic is burned. But the PIC is working !!!!!.
This problem has to be traced. A spare PIC board is found and is ready to switch.

A second inspection of the PIC, what seems burned, it was a glicol drop on the PIC chip and spread on the chip connectors, the drop was dried and the system seems to work ok.

To prevent a few more drops on the electronics the flower pot cover is mounted.

GENERATION OF NEW INTERACTION MATRIX

waiting for the secondary works

-- JuanCarlosGuerra - 2009-09-11
Topic revision: r2 - 11 Sep 2009, JuanCarlosGuerra
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback